OpenStudy (anonymous):

what is 1/(secx+1)+1/(secx-1)=???

sam (.sam.):

Simplify?

sam (.sam.):

(1)/(secx+1)+(1)/(secx-1) (1)/(secx+1) * (secx-1)/(secx-1) + (1)/(secx-1)*(secx+1)/(secx+1) (secx-1)/((secx-1)(secx+1)) + (secx+1)/((secx-1)(secx+1)) ((secx-1)+(secx+1)) / ((secx-1)(secx+1)) 2secx / ((secx-1)(secx+1)) 2secx/(sec^(2)x-1) 2secx / (tan^(2)x) 2secx / ((sin^(2)x)/(cos^(2)x)) (2secxcos^(2)x) / (sin^(2)x) 2cosx/(sin^(2)x) 2cotx/sinx

OpenStudy (mertsj):

\[\frac{\sec x-1+\sec x+1}{\sec ^2x-1}=\frac{2\sec x}{\tan^2x}\]